OpenStudy (shamil98):

What is the value of \[\int\limits_{}^{} \int\limits_{}^{} \int\limits_{R}^{} \cos(x + y + z) ~dx~dy~dz\] \[\ \text{Where R is the region bounded by:}\] \[\ x \ge 0 , y \ge 0 , z \ge 0 ~~\text{and}~~~x+y+z \le 2 \pi \]

OpenStudy (anonymous):

._. i feel so stupid looking at this

OpenStudy (anonymous):



OpenStudy (anonymous):

can you describe the solid?

OpenStudy (shamil98):

That's all the info given.

OpenStudy (shamil98):

looks like a triangle prism

OpenStudy (shamil98):

from your drawing

OpenStudy (shamil98):

oh wait that's a pyramid.

OpenStudy (anonymous):

z ranges from 0 to the plane yes?

OpenStudy (shamil98):

yeah

OpenStudy (anonymous):

what about x and y?

OpenStudy (shamil98):

the same?

OpenStudy (anonymous):



OpenStudy (anonymous):

what is the equation that describe the line?

OpenStudy (shamil98):

\[x + y + z \le 2{\pi}\] ?.. I don't fully understand how to do this ... Could you give an approach without using a graph ? xD

OpenStudy (anonymous):

Well, most of the time, it's impossible to describe a solid without drawing, but in this case, it's possible (I guess because I already had this image in my mind, so I can do it without drawing) the plane intersect the xy-plane at the line x + y = 2pi (just plug 0 in for z)

OpenStudy (shamil98):

ok

OpenStudy (anonymous):

so y = 2pi - x yes?

OpenStudy (shamil98):

Yeah.

OpenStudy (anonymous):

x ranges from 0 to 2pi y ranges from 0 to 2pi - x z ranges from 0 to 2pi - x - y those are the limits of integration. From here on, it's all computation, which I sense it's gonna be quite tedious. If you already how to integrate, just use wolfram alpha

OpenStudy (shamil98):

yeah i know how to integrate from here , thanks for the clarification on the limits

OpenStudy (anonymous):

yw

OpenStudy (shamil98):

I tried putting it in , but I got exceeded time limit.. xD do you have wolframpro?

OpenStudy (anonymous):

no, i just use the website.